What would X= And what would Y= if,
-2x - 3y = -3
and
4x + 5y = 3
I Dont understand how to do this, since you cant use substitution ( i dont think)
And you cant change the signs since they dont have anyn umbers that are alike.
Ive been stuck on this problem for 11minutes now. :-(
i Just need to know what X and Y equal.
And if you could show me How you do it so ill understand for another problem like this i might run into.
Thank you So much,
Keith.

Answer by stanbon(75887)

-2x - 3y = -3
and
4x + 5y = 3
---------------------
Multiply thru the 1st equation by 5.
Multiply thru the 2nd equation by -3.
----------------------------
-10x -15y = -15
-12x -15y = -9
---------------
Subtract the 2nd equation from the 1st to solve for "x":
2x = -6
x = -3
------------
Substitute that into one of the original equations to solve for "y":
-2*-3 -3y = -3
Divide thru by -3 to get:
-2 +y = 1
y = 3
==============
Cheers,
Stan H.